OpenStudy (anonymous):

find the measure and the bases of the median. am gonna draw it :) please helppp me.?

OpenStudy (anonymous):



OpenStudy (anonymous):

What do you know about the length of median in relation to the bases?

OpenStudy (anonymous):

that's why i'm asking for answers :)) i know that bases are the parallel sides and median is the segment joining midpoint of the legs.

OpenStudy (anonymous):

The length of the median is the average length of the bases, so length of median = (length base one + length base two)/2 Can you fill it in?

OpenStudy (anonymous):

oh yeah :) can you solve it? so what's the answer? :D hahaha.

OpenStudy (anonymous):

4x - 3 = (x + 14 + 2x + 10)/2 4x - 3 = (3/2)x + 12 (5/2)x = 15 x = ?

OpenStudy (anonymous):

24? it's wrong right?

OpenStudy (anonymous):

the median?

OpenStudy (anonymous):

and x=8? haha am i wrong or right?

OpenStudy (anonymous):

 x is 6, to get the length of the median and bases just plug in x in the equations :) So the median is 4 * 6 - 3 and the bases are 14 + 6 and 2 * 6 + 10